# ProteinBench: A Holistic Evaluation of Protein Foundation Models

### v1

@@ 1 Introduction
Proteins are fundamental molecules playing pivotal roles in a vast array of biological processes, from enzymatic catalysis and signal transduction to structural support and immune response. Their functions are determined by their amino acid sequences, often mediated through folding into specific three-dimensional structures. Understanding the complex interplay between protein sequence, structure, and function is crucial for advancing science and engineering spanning pharmaceuticals, agriculture, specialty chemicals, and biofuels (Kuhlman & Bradley, 2019).
@@ 3.1.4 Structure and sequence co-design
∗: We have tried our best to reproduce all models according to the instructions in their respective codebases, using publicly available model weights. However, some results may differ from those reported in the original studies. We welcome any feedback and corrections to help us make timely updates in the future.
@@
ProteinGenerator | 0.810.14 | 9.264.13 | 0.22 | 0.71 | 0.690.17 | 17.005.52 | 0.18 | 0.73 |
ProtPardelle* | 0.690.08 | 14.913.45 | 0.04 | 0.72 | 0.440.12 | 43.159.86 | 0.60 | 0.69 |
Multiflow | 0.960.06 | 2.143.24 | 0.58 | 0.71 | 0.950.07 | 2.713.65 | 0.62 | 0.71 |
ESM3* | 0.590.21 | 25.520.68 | 0.52 | 0.73 | 0.640.20 | 26.7221.08 | 0.46 | 0.78 |
@@ 3.2 Protein Conformation Prediction
In the second part of ProteinBench, we focus on conformation prediction, another class of cross-modality tasks aimed at predicting protein structures (conformations) from their sequences. While the current models are based on a body of work distinct from the design tasks, the ability to predict protein conformations provides insight into a model’s understanding of the physics and dynamics of protein structures. This capability is essential for future protein foundation models to fully understand, predict, and design proteins that embody the key sequence-structure-function relationships
The development of conformation prediction models is still in its early stages, with only a handful of exploratory methods proposed. A comprehensive comparison between these methods has yet to be conducted. To the best of our knowledge, this is the first benchmark study on current conformation prediction models that includes the major strategies proposed to date: (1) perturbing the sequence input of folding models (Del Alamo et al., 2022; Wayment-Steele et al., 2024); (2) perturbing protein structures through a structural-only diffusion model (Lu et al., 2024); (3) training generative models on large-scale structural data from experiments or simulations (Jing et al., 2023; 2024; Wang et al., 2024c; Zheng et al., 2024); (4) improving the conformational sampling using physical models (Zheng et al., 2024; Wang et al., 2024c).
@@ 3.2.1 Protein Folding: single-state prediction
While most folding models, such as AlphaFold2 (Jumper et al., 2021) and ESMFold (Lin et al., 2023), are not generative in nature, we still consider them “protein foundation models” for conformation prediction because (1) they are trained on large amounts of structural and/or sequence data; (2) they have played a fundamental role in understanding sequence-structure relationships; and (3) they are closely related to foundation models like ESM2 (Rives et al., 2021) and AlphaFold3 (Abramson et al., 2024), and serve as pivotal building blocks for conformation prediction models (Jing et al., 2024; Wang et al., 2024c). Therefore, it is essential to benchmark their performance in discussions of protein conformation prediction.”
In Table 7, we summarize the results of folding models on CAMEO2022. For AlphaFold2, OpenFold, and EigenFold, we predict five structures and report the best structure based on the model’s internal confidence scores. Consistent with previous reports (Jing et al., 2023), folding models (AlphaFold2, OpenFold, RosettaFold2) based on Multiple Sequence Alignment (MSA) outperform folding models (ESMFold, EigenFold) based on protein language models. AlphaFold2 and its faithful reproduction, OpenFold, demonstrate the best performance across all accuracy metrics. The quality of the predicted structures is comparable among the models (except for EigenFold) with minimal inter-residue clashes or bond breaks. EigenFold (Jing et al., 2023) is one of the first diffusion generative models to claim both protein folding and conformation prediction capabilities; however, its performance may be limited by several design factors: it is built on OmegaFold Wu et al. (2022), uses a coarse-grained representation with only alpha carbons, and has a small model size of 572K trainable parameters. In summary, although the MSA search is time-consuming, AlphaFold2 and OpenFold still achieve the best performance in both accuracy and structural quality for protein folding, providing a strong foundation for protein conformation prediction.
Accuracy | Quality | ||||||
TM-score ↑ | RMSD ↓ | GDT-TS ↑ | lDDT ↑ | CA clash (%) ↓ | CA break (%) ↓ | PepBond break (%) ↓ | |
AlphaFold2 | 0.871/0.952 | 3.21/1.64 | 0.860/0.921 | 0.904/0.933 | 0.3/0.0 | 0.0/0.0 | 4.8/4.1 |
OpenFold | 0.870/0.947 | 3.21/1.59 | 0.856/0.913 | 0.899/0.933 | 0.4/0.0 | 0.0/0.0 | 2.0/1.7 |
RoseTTAFold2 | 0.859/0.941 | 3.52/1.75 | 0.845/0.903 | 0.892/0.926 | 0.3/0.0 | 0.2/0.0 | 5.5/4.0 |
ESMFold | 0.847/0.929 | 3.98/2.10 | 0.826/0.881 | 0.870/0.907 | 0.3/0.0 | 0.0/0.0 | 4.7/3.4 |
EigenFold* | 0.743/0.823 | 7.65/3.73 | 0.703/0.781 | 0.737/0.810 | 8.0/4.6 | 0.5/0.0 | N/A |
@@ 3.2.2 Multiple-state prediction
In this section, we evaluate the performance of predicting multiples conformational states for five conformation prediction models and their variants: EigenFold Jing et al. (2023), MSA-subsampling (Del Alamo et al., 2022), Str2Str (Lu et al., 2024), AlphaFlow/ESMFlow Jing et al. (2024) and ConfDiff (Wang et al., 2024c). Here we highlight the key differences of these methods. MSA-subsampling perturbs OpenFold’s model input by reducing the number of input MSAs (referred to as “depth”), allowing the sampling of different conformations for a protein. Str2Str uses a structure-only diffusion model (i.e., a backbone design model) to generate conformations by perturbing initial folding predictions through a forward-backward diffusion process. The level of perturbation is controlled by the maximum diffusion time, , and ensemble outputs are generated by sampling structures at various diffusion times . EigenFold, AlphaFlow/ESMFlow, and ConfDiff take a similar approach by training diffusion or flow models through fine-tuning folding models using structural data from PDB. While AlphaFlow/ESMFlow open all layers of the original model for fine-tuning, EigenFold and ConfDiff only use pretrained representations from folding models and train a lightweight module for score or structural prediction. Additionally, both AlphaFlow/ESMFlow and ConfDiff provide versions further fine-tuned on a recent MD conformation dataset (Vander Meersche et al., 2024), indicated by the “-MD” suffix. ConfDiff (Wang et al., 2024c) introduced two guidance techniques to improve conformational sampling: (1) classifier-free guidance, which combines a sequence-conditioned conformation model with an unconditional (structure-only) model to explore conformational space (ConfDiff-ClsFree), and (2) energy/force guidance, which directs sampling toward regions with lower potential energy (ConfDiff-Energy/Force) through auxiliary prediction modules for intermediate energy/force guidance. However, such physical prediction modules are dataset-specific and are only available for the BPTI dataset.
RMSDens ↓ | RMSD Cluster 3 ↓ | Diversity | Quality | |||||||||||||||||
N=10 | N=100 | N=500 | N=1000 | N=10 | N=100 | N=500 | N=1000 |
EigenFold | 1.560.02 | 1.500.01 | 1.470.01 | 1.460.00 | 2.540.03 | 2.480.01 | 2.460.01 | 2.460.01 | 0.85 | 1.4 | 4.3 | N/A | ||||||||
MSA-depth256 | 1.570.01 | 1.540.01 | 1.520.00 | 1.520.01 | 2.510.02 | 2.470.02 | 2.450.02 | 2.450.02 | 0.20 | 0.0 | 0.0 | 9.2 | ||||||||
MSA-depth64 | 1.600.02 | 1.540.01 | 1.510.01 | 1.500.01 | 2.480.03 | 2.400.04 | 2.350.04 | 2.330.03 | 0.55 | 0.0 | 0.0 | 7.9 | ||||||||
MSA-depth32 | 1.670.05 | 1.530.04 | 1.450.04 | 1.410.02 | 2.390.15 | 2.210.15 | 1.930.15 | 1.870.06 | 2.14 | 0.6 | 0.0 | 10.6 | ||||||||
Str2Str-ODE () | 2.360.10 | 2.190.06 | 2.100.02 | 2.080.01 | 3.030.17 | 2.680.12 | 2.600.05 | 2.560.02 | 1.86 | 0.0 | 0.0 | 13.9 | ||||||||
Str2Str-SDE () | 2.830.23 | 2.480.11 | 2.280.04 | 2.250.03 | 3.420.32 | 2.920.28 | 2.520.14 | 2.480.13 | 3.60 | 0.3 | 0.0 | 16.0 | ||||||||
AlphaFlow-PDB | 1.530.02 | 1.450.01 | 1.420.01 | 1.410.01 | 2.480.04 | 2.430.02 | 2.410.02 | 2.400.01 | 0.86 | 0.0 | 0.0 | 13.2 | ||||||||
AlphaFlow-MD | 1.740.09 | 1.510.04 | 1.450.02 | 1.430.02 | 2.440.06 | 2.320.06 | 2.280.04 | 2.240.00 | 1.26 | 0.0 | 0.1 | 26.2 | ||||||||
ESMFlow-PDB | 1.610.04 | 1.490.02 | 1.440.01 | 1.420.01 | 2.470.05 | 2.410.03 | 2.370.03 | 2.350.01 | 0.74 | 0.0 | 0.0 | 6.0 | ||||||||
ESMFlow-MD | 1.660.07 | 1.500.04 | 1.410.03 | 1.400.02 | 2.490.06 | 2.290.09 | 2.200.04 | 2.180.03 | 1.17 | 0.0 | 0.0 | 14.3 | ||||||||
ConfDiff-Open-ClsFree | 1.650.05 | 1.480.05 | 1.410.04 | 1.370.03 | 2.560.05 | 2.300.23 | 2.160.20 | 2.030.13 | 1.77 | 0.5 | 0.0 | 5.5 | ||||||||
ConfDiff-Open-MD | 1.640.04 | 1.500.02 | 1.440.02 | 1.420.02 | 2.490.08 | 2.390.05 | 2.320.03 | 2.310.02 | 1.37 | 0.2 | 0.0 | 4.6 | ||||||||
ConfDiff-ESM-ClsFree | 1.580.05 | 1.450.02 | 1.410.01 | 1.390.00 | 2.500.05 | 2.390.03 | 2.350.03 | 2.330.02 | 1.52 | 0.5 | 0.0 | 7.5 | ||||||||
ConfDiff-ESM-MD | 1.610.03 | 1.470.02 | 1.420.01 | 1.400.01 | 2.450.08 | 2.320.06 | 2.260.04 | 2.240.01 | 1.42 | 0.1 | 0.0 | 5.0 | ||||||||
ConfDiff-ESM-Energy | 1.630.06 | 1.470.01 | 1.430.01 | 1.420.01 | 2.550.07 | 2.430.04 | 2.410.02 | 2.400.01 | 1.26 | 0.1 | 0.0 | 7.5 | ||||||||
ConfDiff-ESM-Force | 1.580.06 | 1.440.03 | 1.370.02 | 1.360.01 | 2.450.09 | 2.330.07 | 2.230.06 | 2.220.06 | 1.76 | 0.1 | 0.0 | 8.9 |
In Table 8, we summarize the results on predicting the five structural clusters of BPTI. Specifically, the ensemble accuracy of five clusters (RMSDens) and the accuracy of Cluster 3 (RMSD Cluster 3), the most difficult to sample, are evaluated by bootstrapping at different sample sizes. ConfDiff models with classifier-free guidance demonstrated the best performance in overall accuracy (RMSDens) across most sample sizes, highlighting its greater effectiveness on BPTI compared to fine-tuning on MD conformation data. Consistent with Wang et al. (2024c), ConfDiff-ESM-Force achieved the highest ensemble accuracy, suggesting that incorporating physical information can improve the sampling of high-accuracy conformations. For the task of sampling Cluster 3, MSA subsampling, despite being a simple approach, was able to generate conformations most likely to capture this remote state. As the MSA depth decreases, sample diversity increases, allowing the model to sample more closely to Cluster 3. In comparison, Str2Str models do not perform well on this task, potentially because the structure-only approach does not ensure that the perturbed structure remains faithful to the provided sequence, leading to overall poorer performance. EigenFold also shows lower diversity, which may limit its effectiveness in sampling diverse conformations. AlphaFlow/ESMFlow also demonstrated competitive performance. Fine-tuning on the MD dataset provided higher diversity and improved accuracy in sampling Cluster 3 compared to their PDB-trained base models. However, we also observed a more pronounced decline in quality due to fine-tuning on MD conformation data, specifically with an increased rate of peptide bond breaking between residues.
Accuracy | Diversity | Quality | ||||||
apo-TM ↑ | holo-TM ↑ | TMens ↑ | Pairwise TM | CA clash % ↓ | CA break% ↓ |
apo model | 1.000/1.000 | 0.790/0.821 | 0.895/0.910 | N/A | N/A | N/A | N/A | |
EigenFold | 0.831/0.862 | 0.864/0.900 | 0.847/0.874 | 0.907/0.958 | 3.6/1.2 | 0.3/0.1 | N/A | |
MSA-depth256 | 0.845/0.882 | 0.889/0.936 | 0.867/0.894 | 0.978/0.993 | 0.2/0.0 | 0.0/0.0 | 4.6/4.0 | |
MSA-depth64 | 0.844/0.877 | 0.883/0.927 | 0.863/0.906 | 0.950/0.980 | 0.2/0.0 | 0.0/0.0 | 5.7/5.0 | |
MSA-depth32 | 0.824/0.865 | 0.857/0.905 | 0.841/0.882 | 0.864/0.924 | 0.2/0.0 | 0.0/0.0 | 8.9/7.3 | |
Str2Str-ODE () | 0.762/0.791 | 0.778/0.816 | 0.770/0.794 | 0.954/0.956 | 0.2/0.0 | 0.0/0.0 | 14.0/12.4 | |
Str2Str-ODE () | 0.766/0.797 | 0.781/0.818 | 0.774/0.797 | 0.872/0.871 | 0.2/0.0 | 0.0/0.0 | 14.7/12.9 | |
Str2Str-SDE () | 0.682/0.703 | 0.693/0.717 | 0.688/0.712 | 0.760/0.748 | 0.2/0.1 | 1.5/1.5 | 22.6/21.3 | |
Str2Str-SDE () | 0.680/0.685 | 0.689/0.718 | 0.684/0.697 | 0.639/0.604 | 0.2/0.1 | 1.4/1.4 | 21.1/19.6 | |
AlphaFlow-PDB | 0.855/0.896 | 0.891/0.942 | 0.873/0.900 | 0.924/0.955 | 0.3/0.0 | 0.0/0.0 | 6.6/6.7 | |
AlphaFlow-MD | 0.857/0.888 | 0.863/0.913 | 0.860/0.892 | 0.894/0.918 | 0.2/0.0 | 0.0/0.0 | 20.8/20.6 | |
ESMFlow-PDB | 0.849/0.878 | 0.882/0.924 | 0.866/0.900 | 0.935/0.952 | 0.3/0.1 | 0.0/0.0 | 4.8/4.7 | |
ESMFlow-MD | 0.851/0.882 | 0.864/0.908 | 0.858/0.890 | 0.897/0.922 | 0.1/0.0 | 0.0/0.0 | 10.9/10.9 | |
ConfDiff-Open-ClsFree | 0.838/0.886 | 0.879/0.927 | 0.859/0.885 | 0.870/0.898 | 0.8/0.6 | 0.0/0.0 | 5.8/5.6 | |
ConfDiff-Open-MD | 0.839/0.881 | 0.874/0.918 | 0.857/0.890 | 0.863/0.892 | 0.4/0.2 | 0.0/0.0 | 6.8/6.8 | |
ConfDiff-ESM-ClsFree | 0.837/0.883 | 0.864/0.907 | 0.850/0.887 | 0.846/0.869 | 0.7/0.6 | 0.0/0.0 | 4.6/4.5 | |
ConfDiff-ESM-MD | 0.836/0.877 | 0.862/0.908 | 0.849/0.892 | 0.846/0.875 | 0.3/0.2 | 0.0/0.0 | 4.1/4.0 |
We next move on to a larger dataset, apo, which includes 91 proteins with ligand-binding-induced conformation changes (Table 9). In this task, models are required to predict both the unbound (apo) and bound (holo) structures. Interestingly, we found that the best-performing models are those that most closely resemble folding models (e.g., MSA-depth256, AlphaFlow-PDB). Despite using a small perturbation level (), Str2Str does not accurately predict either the apo or holo structures. Strategies to improve sample diversity, such as decreasing MSA depth, fine-tuning on MD conformation data, or using classifier-free guidance, generally do not improve (and sometimes even harm) the TMens score. Additionally, we included a baseline model that always predicts the perfect apo structure (the apo model), which achieved a higher TMens score than current models. These findings suggest that a strong folding model improves sampling quality, but current performance struggles to outperform a “perfect” folding model, as the apo-holo conformation challenge remains difficult for existing protein conformation models.
In the multiple-state prediction task, we observed that strategies such as MSA subsampling, classifier-free guidance, force guidance, and training on MD conformation data improved sample diversity and accuracy on the BPTI dataset. However, most of these strategies failed to improve the dual-state prediction task for apo-holo conformational changes
@@ 3.2.3 Distribution prediction
In this final task, we benchmark models on the ATLAS test set, which includes 82 proteins, and focus on each model’s ability to recover the conformational distribution observed in classic protein molecular dynamics simulations. The results are summarized in Table 10. For comparison, we include reference performances of (1) i.i.d. samples (MD iid) from MD-generated structures and (2) 250 consecutive samples, corresponding to 2.5 ns of simulation time (MD 2.5 ns).
Overall, generative models trained to sample protein conformations from sequence (AlphaFlow/ESMFlow, ConfDiff) perform considerably better than perturbation-based methods (MSA subsampling and Str2Str) across nearly all accuracy metrics of flexibility prediction, distributional accuracy, and ensemble observables. In our experiments, adjusting perturbation levels for Str2Str (the maximum forward time ) and MSA-subsampling (the MSA depth) did not improve distribution-related metrics, suggesting that perturbation alone may not be sufficient to accurately capture the sample distribution. We identified two factors consistently improve the model performance of AlphaFlow/ESMFlow and ConfDiff: (1) choosing a strong base folding model (e.g., AlphaFold or OpenFold), which, while potentially reducing sample diversity, improves distribution-related metrics; and (2) fine-tuning process on MD conformation data, which further enhances the models’ ability to predict the target distribution. The latter again emphasizes the importance of aligning the model’s distribution to the target distribution through supervised methods, rather than relying solely on conformation exploration strategies (e.g., classifier-free guidance), to accurately predict distributions. In addition, the results align with previous tasks regarding the trade-offs between diversity, prediction performance, and sample quality: for example, while fine-tuning on MD conformation data improves sample diversity and prediction performance for AlphaFlow/ESMFlow, it also significantly increases the level of peptide bond breaking in these models.
While current conformation prediction models have shown promising signs in capturing dynamics-related features and approximating conformation distributions, it is important to note that there remains a clear gap between the performance of these models and that of classic MD simulations, even at short simulation times (e.g., 2.5 ns). Achieving performance comparable to i.i.d. sampling from MD conformational samples has yet to be achieved.
@@
MD iid | 2.76 | 1.63 | 0.96 | 0.97 | 0.99 | 0.71 | 0.76 | 0.70 | 93.9 | |||||||||||||||
MD 2.5 ns | 1.54 | 0.98 | 0.89 | 0.85 | 0.85 | 2.21 | 1.57 | 1.93 | 36.6 | |||||||||||||||
EigenFold | 5.96 | N/A | -0.04 | N/A | N/A | N/A | 2.35 | 7.96 | 12.2 | |||||||||||||||
MSA-depth256 | 0.84 | 0.53 | 0.25 | 0.34 | 0.59 | 3.63 | 1.83 | 2.90 | 29.3 | |||||||||||||||
MSA-depth64 | 2.03 | 1.51 | 0.24 | 0.30 | 0.57 | 4.00 | 1.87 | 3.32 | 18.3 | |||||||||||||||
MSA-depth32 | 5.71 | 7.96 | 0.07 | 0.17 | 0.53 | 6.12 | 2.50 | 5.67 | 17.1 | |||||||||||||||
Str2Str-ODE (t=0.1) | 1.66 | N/A | 0.13 | N/A | N/A | N/A | 2.12 | 4.42 | 6.1 | |||||||||||||||
Str2Str-ODE (t=0.3) | 3.15 | N/A | 0.12 | N/A | N/A | N/A | 2.23 | 4.75 | 9.8 | |||||||||||||||
Str2Str-SDE (t=0.1) | 4.74 | N/A | 0.10 | N/A | N/A | N/A | 2.54 | 8.84 | 9.8 | |||||||||||||||
Str2Str-SDE (t=0.3) | 7.54 | N/A | 0.00 | N/A | N/A | N/A | 3.29 | 12.28 | 7.3 | |||||||||||||||
AlphaFlow-PDB | 2.58 | 1.20 | 0.27 | 0.46 | 0.81 | 2.96 | 1.66 | 2.60 | 37.8 | |||||||||||||||
AlphaFlow-MD | 2.88 | 1.63 | 0.53 | 0.66 | 0.85 | 2.68 | 1.53 | 2.28 | 39.0 | |||||||||||||||
ESMFlow-PDB | 3.00 | 1.68 | 0.14 | 0.27 | 0.71 | 4.20 | 1.77 | 3.54 | 28.0 | |||||||||||||||
ESMFlow-MD | 3.34 | 2.13 | 0.19 | 0.30 | 0.76 | 3.63 | 1.54 | 3.15 | 25.6 | |||||||||||||||
ConfDiff-Open-ClsFree | 3.68 | 2.12 | 0.40 | 0.54 | 0.83 | 2.92 | 1.50 | 2.54 | 46.3 | |||||||||||||||
ConfDiff-Open-PDB | 2.90 | 1.43 | 0.38 | 0.51 | 0.82 | 2.97 | 1.57 | 2.51 | 34.1 | |||||||||||||||
ConfDiff-Open-MD | 3.43 | 2.21 | 0.59 | 0.67 | 0.85 | 2.76 | 1.44 | 2.25 | 35.4 | |||||||||||||||
ConfDiff-ESM-ClsFree | 4.04 | 2.84 | 0.31 | 0.43 | 0.82 | 3.82 | 1.72 | 3.06 | 37.8 | |||||||||||||||
ConfDiff-ESM-PDB | 3.42 | 2.06 | 0.29 | 0.40 | 0.80 | 3.67 | 1.70 | 3.17 | 34.1 | |||||||||||||||
ConfDiff-ESM-MD | 3.91 | 2.79 | 0.35 | 0.48 | 0.82 | 3.67 | 1.66 | 2.89 | 39.0 | |||||||||||||||

### v2

@@ 1 Introduction
Proteins are fundamental molecules playing pivotal roles in a vast array of biological processes, from enzymatic catalysis and signal transduction to structural support and immune response. Their amino acid sequences determine their functions, often mediated through folding into specific three-dimensional structures. Understanding the complex interplay between protein sequence, structure, and function is crucial for advancing science and engineering spanning pharmaceuticals, agriculture, specialty chemicals, and biofuels (Kuhlman & Bradley, 2019).
@@ 3.1.4 Structure and sequence co-design
ProteinGenerator | 0.810.14 | 9.264.13 | 0.22 | 0.71 | 0.410.19 | 33.9115.10 | 0.18 | 0.73 |
ProtPardelle | 0.940.02 | 2.070.73 | 0.05 | 0.73 | 0.410.10 | 41.108.85 | 0.14 | 0.65 |
Multiflow | 0.960.06 | 2.143.24 | 0.58 | 0.71 | 0.830.15 | 8.487.02 | 0.67 | 0.68 |
ESM3* | 0.590.21 | 25.520.68 | 0.52 | 0.73 | 0.540.20 | 33.7021.08 | 0.37 | 0.77 |
@@ 3.2 Protein Conformation Prediction
Protein conformation prediction infers the 3D conformations of proteins from their sequences, evaluating models based on their understanding of structure, dynamics, and ultimately functions. We begin by benchmarking common folding models, as they play a crucial role in the development of conformation prediction models. We then compare five recent studies that explore difference strategies to extend folding to conformation prediction, focusing on multiple-state and distribution prediction tasks. Key differences between these approaches are outlined in Appendix B.3.4.
@@ 3.2.1 Protein Folding: single-state prediction
As shown in Table 7, Multiple Sequence Alignment (MSA) based folding models (AlphaFold2, OpenFold, RoseTTAFold2) outperforms protein language model-based approaches (ESMFold, EigenFold). While the predicted structure quality is comparable among the all-atom resolution models, AlphaFold2 and its reproduction, OpenFold, achieve the best performance across all accuracy metrics, offering a strong foundation for conformation prediction.
Accuracy | Quality | |||||
TM-score ↑ | RMSD ↓ | GDT-TS ↑ | lDDT ↑ | CA clash (%) ↓ | PepBond break (%) ↓ | |
AlphaFold2 | 0.871/0.952 | 3.21/1.64 | 0.860/0.921 | 0.904/0.933 | 0.3/0.0 | 4.8/4.1 |
OpenFold | 0.870/0.947 | 3.21/1.59 | 0.856/0.913 | 0.899/0.933 | 0.4/0.0 | 2.0/1.7 |
RoseTTAFold2 | 0.859/0.941 | 3.52/1.75 | 0.845/0.903 | 0.892/0.926 | 0.3/0.0 | 5.5/4.0 |
ESMFold | 0.847/0.929 | 3.98/2.10 | 0.826/0.881 | 0.870/0.907 | 0.3/0.0 | 4.7/3.4 |
EigenFold | 0.743/0.823 | 7.65/3.73 | 0.703/0.781 | 0.737/0.810 | 8.0/4.6 | N/A |
@@ 3.2.2 Multiple-state prediction
We first investigate the results on BPTI with the best model of each study highlighted in Table 8 and the complete evaluation in Table 13. The classifier-free guidance in ConfDiff (Wang et al., 2024c) achieved performance comparable to fine-tuning on MD conformation data and, when combined with force guidance, delivered the best overall accuracy (RMSDens). This suggests that incorporating structural exploration and physical constraints can enhance the sampling of high-accuracy conformations. However, structural exploration alone may be error-prone as Str2Str (Lu et al., 2024), the structure-only models, showed low accuracy even with small perturbations. Other strategies to enhance conformation sampling also showed improved performance: MSA subsampling (Del Alamo et al., 2022) with reduced MSA depth excelled at sampling Cluster 3, the most difficult state to capture, and ESMFlow (Jing et al., 2024) fine-tuned on the MD dataset showed improved diversity and accuracy compared to the PDB-trained base model. However, these approaches also experienced a decline in quality, with increased CA clashing or peptide bond breaking. Lastly, for most methods, increasing the sample depth () significantly improved expected accuracy, suggesting that a sufficient sample size is essential for thorough evaluation.
RMSDens (Å) ↓ | RMSD (Å) Cluster 3 ↓ | Diversity | Quality | ||||||||
N=10 | N=100 | N=1000 | N=10 | N=100 | N=1000 |
CA clash% ↓ | PepBond break%↓ | |||
EigenFold | 1.560.02 | 1.500.01 | 1.460.00 | 2.540.03 | 2.480.01 | 2.460.01 | 0.85 | 1.4 | N/A | ||
MSA-depth32 | 1.660.03 | 1.540.04 | 1.410.02 | 2.430.06 | 2.190.16 | 1.850.05 | 2.14 | 0.6 | 10.6 | ||
Str2Str-ODE () | 2.400.12 | 2.200.05 | 2.090.01 | 3.000.20 | 2.730.12 | 2.580.05 | 1.86 | 0.0 | 13.9 | ||
ESMFlow-MD | 1.680.06 | 1.470.04 | 1.390.03 | 2.440.11 | 2.270.10 | 2.180.02 | 1.17 | 0.0 | 14.3 | ||
ConfDiff-ESM-Force | 1.580.04 | 1.430.03 | 1.360.01 | 2.440.06 | 2.350.05 | 2.240.06 | 1.76 | 0.1 | 8.9 |
Accuracy | Diversity | Quality | |||||
apo-TM ↑ | holo-TM ↑ | TMens ↑ | Pairwise TM | CA clash % ↓ |
apo model | 1.000 | 0.790 | 0.895 | N/A | N/A | N/A | |
EigenFold | 0.831 | 0.864 | 0.847 | 0.907 | 3.6 | N/A | |
MSA-depth256 | 0.845 | 0.889 | 0.867 | 0.978 | 0.2 | 4.6 | |
Str2Str-ODE () | 0.766 | 0.781 | 0.774 | 0.872 | 0.2 | 14.7 | |
AlphaFlow-PDB | 0.855 | 0.891 | 0.873 | 0.924 | 0.3 | 6.6 | |
ConfDiff-Open-PDB | 0.847 | 0.886 | 0.867 | 0.909 | 0.5 | 5.5 |
apo-holo is a more challenging dataset where models are required to predict both the unbound (apo) and bound (holo) conformations induced by ligand binding. As shown in Table 9 and Table 14, strategies to enhance conformation diversity did not improve the TMens score, and the best-performing models closely resemble folding models. In comparison, a baseline model that consistently predicts the perfect apo structure achieved a higher TMens score. These findings suggest that these models struggle to accurately sample apo-holo conformation changes, and higher accuracy may stem from using a stronger folding model.
In summary, strategies like MSA subsampling, guidance during diffusion, or training on MD conformation data can improve sample diversity and enhance ensemble accuracy for certain local dynamics (as in BPTI). However, they may not be sufficient to capture the complex dynamics involved in processes like apo-holo conformational changes.
@@ 3.2.3 Distribution prediction
In this section, we focus on predicting the conformational distributions observed in classic molecular dynamics simulations. The results are summarized in Table 10 and Table 15. We include reference performances of (1) MD iid: i.i.d. samples from reference MD trajectories and (2) MD 2.5 ns: consecutive samples from the trajectories corresponding to 2.5 ns of simulation.
Overall, generative models trained to sample protein conformations (AlphaFlow, ConfDiff) significantly outperform perturbation-based methods (MSA subsampling and Str2Str), regardless of perturbation levels. Using a strong folding model like AlphaFold2 generally results in higher accuracy. Classifier-free guidance in ConfDiff improved distribution sampling but were less effective than direct fine-tuing on MD data, highlighting the importance of large-scale conformational data for future models. Additionally, the trade-offs between diversity, prediction performance, and sample quality persist: fine-tuning on MD data improves sample diversity and performance for AlphaFlow but decreases sample quality.
AlphaFlow and ConfDiff models fine-tuned on MD data have shown promise in capturing the conformational distributions, achieving performance comparable to that of 2.5 ns MD simulations on some metrics. However, a key goal for these models is to achieve i.i.d. sampling equivalent to long-term MD simulations, and benchmark results reveal a remaining gap in reaching this objective.
@@
MD iid | 2.76 | 1.63 | 0.96 | 0.97 | 0.99 | 0.67 | 0.73 | 0.71 | 93.9 | |||||||||||||||
MD 2.5ns | 1.54 | 0.98 | 0.89 | 0.85 | 0.85 | 2.22 | 1.55 | 1.89 | 36.6 | |||||||||||||||
EigenFold | 5.96 | N/A | -0.03 | N/A | N/A | N/A | 2.31 | 7.96 | 12.2 | |||||||||||||||
MSA-depth256 | 0.83 | 0.53 | 0.25 | 0.34 | 0.59 | 3.60 | 1.79 | 2.91 | 29.3 | |||||||||||||||
Str2Str-ODE () | 1.66 | N/A | 0.13 | N/A | N/A | N/A | 2.14 | 4.39 | 6.1 | |||||||||||||||
AlphaFlow-MD | 2.87 | 1.63 | 0.53 | 0.66 | 0.85 | 2.64 | 1.55 | 2.29 | 39.0 | |||||||||||||||
ConfDiff-Open-MD | 3.43 | 2.21 | 0.59 | 0.67 | 0.85 | 2.75 | 1.41 | 2.27 | 35.4 | |||||||||||||||